A car leaves Canberra heading towards Sydney on the freeway at 5pm travelling at a constant speed of 80km/hr. forty-five minutes later a second car leaves in pursuit of the fist car travelling at a constant speed of 110 km/hr. When will the second car catch up to the second car?
Answer by mananth(16949)

A car leaves Canberra heading towards Sydney on the freeway at 5pm travelling at a constant speed of 80km/hr. forty-five minutes later a second car leaves in pursuit of the fist car travelling at a constant speed of 110 km/hr. When will the second car catch up to the second car?
let second car take x hours to catch up
speed = 110 km/h
distance = 110 x
First car
speed = 80 km/h
time = x+3/4 hour
Distance = 80(x+ 3/4)
110x= 80(x+3/4)
11/8 x = x+3/4
11/8 x -x= 3/4
3x/8 = 3/4
3x=6
x=2 hours
5.45+2 hours
7.45 pm